# across languages

### v1

@@ Appendix C Model Outputs
{CJK}UTF8minXだと思う。 Japanese weak
{CJK}UTF8minXが正しいと思う。 Japanese weak
{CJK}UTF8minXが正解だと思う。 Japanese weak
{CJK}UTF8minXだろう。 Japanese weak
{CJK}UTF8minおそらくXだと思う。 Japanese weak
{CJK}UTF8minXだと思われる。 Japanese moderate
{CJK}UTF8minXが正しいと思われる。 Japanese moderate
{CJK}UTF8minXだと考えられる。 Japanese moderate
{CJK}UTF8minXである可能性が高い。 Japanese moderate
{CJK}UTF8minXで間違いないと思う。 Japanese moderate
{CJK}UTF8minXだと確信している。 Japanese strong
{CJK}UTF8minXに違いない。 Japanese strong
{CJK}UTF8minXで間違いない。 Japanese strong
{CJK}UTF8minXに間違いない。 Japanese strong
{CJK}UTF8min確実にXだ。 Japanese strong
{CJK}UTF8min我觉得这可能是 X Mandarin weak
{CJK}UTF8min我猜它是 X Mandarin weak
{CJK}UTF8min它可能是 X Mandarin weak
{CJK}UTF8min我不太确定,但可能是 X Mandarin weak
{CJK}UTF8min我不确定,但它有可能是 X Mandarin weak
{CJK}UTF8min我认为这是合理的选择,它是 X Mandarin moderate
{CJK}UTF8min我肯定它是 X Mandarin moderate
{CJK}UTF8min我确认它是 X Mandarin moderate
{CJK}UTF8min我认为最有可能的答案是 X Mandarin moderate
{CJK}UTF8min我的直觉告诉我,它是 X Mandarin moderate
{CJK}UTF8min我对这个答案非常确定,它是 X Mandarin strong
{CJK}UTF8min我绝对认为它是 X Mandarin strong
{CJK}UTF8min我有很大把握,它是 X Mandarin strong
{CJK}UTF8min我对这个答案有很大的把握,它是 X Mandarin strong
{CJK}UTF8min我有把握认为它是 X Mandarin strong

### v2

@@ Appendix C Model Outputs
Xだと思う。 Japanese weak
Xが正しいと思う。 Japanese weak
Xが正解だと思う。 Japanese weak
Xだろう。 Japanese weak
おそらくXだと思う。 Japanese weak
Xだと思われる。 Japanese moderate
Xが正しいと思われる。 Japanese moderate
Xだと考えられる。 Japanese moderate
Xである可能性が高い。 Japanese moderate
Xで間違いないと思う。 Japanese moderate
Xだと確信している。 Japanese strong
Xに違いない。 Japanese strong
Xで間違いない。 Japanese strong
Xに間違いない。 Japanese strong
確実にXだ。 Japanese strong
我觉得这可能是 X Mandarin weak
我猜它是 X Mandarin weak
它可能是 X Mandarin weak
我不太确定,但可能是 X Mandarin weak
我不确定,但它有可能是 X Mandarin weak
我认为这是合理的选择,它是 X Mandarin moderate
我肯定它是 X Mandarin moderate
我确认它是 X Mandarin moderate
我认为最有可能的答案是 X Mandarin moderate
我的直觉告诉我,它是 X Mandarin moderate
我对这个答案非常确定,它是 X Mandarin strong
我绝对认为它是 X Mandarin strong
我有很大把握,它是 X Mandarin strong
我对这个答案有很大的把握,它是 X Mandarin strong
我有把握认为它是 X Mandarin strong